Rewrite the equation given in Problem 18 to be a function of x. As it turns out, this is very much like Part (b) of Example 4 on p. 656!
18) -3x + 4y = 11

Answer by ankor@dixie-net.com(22740)

Rewrite the equation given in Problem 18 to be a function of x. As it turns out, this is very much like Part (b) of Example 4 on p. 656!
18) -3x + 4y = 11
:
4y = 3x + 11; added 3x to both sides
y = (3/4)x + (11/4); divided both sides by 4
:
y = f(x)
Therefore:
f(x) = (3/4)x + (11/4)